In a small country town, the demand for bread is Pd = 200 - 0.1Q. The supply of bread is Ps = - 50 + 0.4Q. P is the price of bread (in cents) and Q is measured in number of loaves per day
1
Expert's answer
2016-12-14T11:33:12-0500
Pd = Ps
200 – 0.1Q = -50 + 0.4Q
0.5Q = 250
Q = 500 (loaves of bread)
If Q = 500, then Pd = 200-0.1*500 = 150 (cents)
So, we have such results:
Q = 500 and Pd = Ps = 150.